# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 2.1.1. Affine Weyl group
Let $W_{\mathrm{af}}=W\ltimes{\mathrm{Q}}$ be the affine Weyl group. Let $W_{\mathrm{ex}}=W\ltimes\Lambda$ be the extended affine Weyl group. We have $W_{\mathrm{ex}}\simeq W_{\mathrm{af}}\rtimes\Lambda/{\mathrm{Q}}$, where $\Lambda/{\mathrm{Q}}$-acts by automorphisms of the affine root system. Let ${\mathrm{I}}_{\mathrm{af}}$ be the subset of simple reflections in $W_{\mathrm{af}}$. Denote by $t(\lambda)$ the translation by $\lambda\in\Lambda$ in $W_{\mathrm{ex}}$. For any $n\in{\mathbb{Z}}$, there is an $n$-dilated $(-\rho)$-shifted action of $W_{\mathrm{ex}}$ on $\Lambda$ given by
@@
[…]
@@ 2.2. Quantum groups
We set ${\mathfrak{U}}^{t}_{\zeta}$ be the algebra with triangular decomposition ${\mathfrak{U}}^{t}_{\zeta}=u_{\zeta}^{-}\otimes{\mathfrak{U}}_{\zeta}^{0}\otimes u_{\zeta}^{+}$.
@@ 2.3. Module categories
The $\Lambda$-action on $\mathscr{U}_{q}^{0}$ preserves the integral forms ${\mathfrak{U}}^{0}_{q}$ and $U^{0}_{q}$, and specializes to an action on ${\mathfrak{U}}^{0}_{\zeta}$ and $U^{0}_{\zeta}$. Let $U^{0}$ be a Noetherian sub-quotient algebra of ${\mathfrak{U}}_{\zeta}^{0}$ or $U_{\zeta}^{0}$ such that the $\Lambda$-action induces an action on $U^{0}$. Let $U=\bigoplus_{\lambda\in{\mathrm{Q}}}U_{\lambda}$ be a ${\mathrm{Q}}$-graded ${\mathbb{C}}$-algebra with triangular decomposition $U=U^{-}\otimes U^{0}\otimes U^{+}$, such that
@@
and satisfies further conditions as in [37, §2.3]. We abbreviate $U^{\leq}=U^{-}U^{0}$ and $U^{\geq}=U^{+}U^{0}$. A deformation ring $R$ for $U$ is a commutative Noetherian $U^{0}$-algebra. Let $\iota:U^{0}\rightarrow R$ be the structure map. If $U^{0}$ is an augmented algebra, then we will view ${\mathbb{C}}$ as a deformation ring via the counit map $U^{0}\rightarrow{\mathbb{C}}$.
@@
Let $U\text{-}\mathrm{mod}_{R}^{\Lambda}$ be the full subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ consisting of finitely generated $U\otimes R$-modules whose weight spaces are finitely generated $R$-modules. Define the category ${\mathcal{O}}$ for $U$ to be the full subcategory ${\mathcal{O}}^{U}_{R}$ of $U\text{-}\mathrm{mod}_{R}^{\Lambda}$ of modules that are locally unipotent for the action of $U^{+}$. It is an abelian subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
@@
where $R_{\lambda}$ is a $U^{\geq}$-module via $U^{\geq}\twoheadrightarrow U^{0}\xrightarrow{\iota\circ\tau_{\lambda}}R_{\lambda}$. We have an isomorphism $M^{U}(\lambda)_{R}=U^{-}\otimes R_{\lambda}$ as $U^{\leq}$-modules. A Verma flag of an object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ is a (decreasing, separated, ${\mathbb{Z}}_{\leq 0}$-) filtration with composition factors by Verma modules. The composition factors of such an object are called Verma factors, which are independent on the choice of Verma flags.
@@ 2.3.1. Truncated categories
For any $\nu\in\Lambda$, there is a truncated subcategory $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ consists of the module $M$ in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ such that $M_{\mu}=0$ unless $\mu\leq\nu$. Note that $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is a Serre subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$. The category ${\mathcal{O}}^{U,\leq\nu}_{R}:={\mathcal{O}}^{U}_{R}\cap U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ always admits enough projective objects, in contrast to ${\mathcal{O}}^{U}_{R}$. And any object in ${\mathcal{O}}^{U}_{R}$ admits a finite decomposition with each direct factor contained in ${\mathcal{O}}^{U,\leq\nu}_{R}$ for some $\nu\in\Lambda$. Any projective module in ${\mathcal{O}}^{U,\leq\nu}_{R}$ admits Verma flags. If $R$ is a complete local domain with residue field ${\mathbb{F}}$, we write $Q^{U}(\lambda)_{R}^{\leq\nu}$ as the projective cover of $L^{U}(\lambda)_{\mathbb{F}}$ in ${\mathcal{O}}^{U,\leq\nu}_{R}$.
The truncation functor
$U\text{-}\mathrm{Mod}_{R}^{\Lambda}\rightarrow U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu},\quad M\mapsto\ M^{\leq\nu}:=M\big{/}U.\big{(}\bigoplus_{\mu\nleq\nu}M_{\mu}\big{)}$
by taking the maximal quotient in $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is left adjoint to the natural inclusion $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}\hookrightarrow U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
@@
  1. (1)
    We have
    $\mathrm{Hom}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda\leq\mu,$
    and
    $\mathrm{Ext}^{1}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda<\mu.$
  2. (2)
    For any object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ admitting Verma flags, its truncation in $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is the quotient by the submodule extended by its Verma factors in $\{M^{U}(\lambda)_{R}\}_{\lambda\nleq\nu}$.
@@ 2.3.2. Projectives and simples in ${\mathcal{O}}_{R}$
We will drop the superscripts $U$ from the notations above when $U=U^{\mathrm{hb}}_{\zeta}$. We set
@@
Then $S$ is naturally an algebra of ${\mathfrak{U}}^{0}_{\zeta}={\mathbb{C}}\langle K_{\lambda}\rangle_{\lambda\in\Lambda}={\mathbb{C}}[T]$. Note that $S$ is a deformation ring for both ${\mathfrak{U}}^{b}_{\zeta}$ and $U^{\mathrm{hb}}_{\zeta}$.
Lemma 2.2 ([37, Lem 3.1]).
@@
    The projective cover for $L(\lambda)_{\mathbb{C}}$ in ${\mathcal{O}}^{\leq\nu}_{S}$ (resp. ${\mathcal{O}}^{\leq\nu}_{{\mathbb{C}}}$) is $Q(\lambda)_{S}^{\leq\nu}=\tau^{\leq\nu}Q(\lambda)_{S}$ (resp. $Q(\lambda)_{\mathbb{C}}^{\leq\nu}=\tau^{\leq\nu}Q(\lambda)_{\mathbb{C}}$).
@@ 2.3.3. Block decomposition
Till the end of this section, we let $U$ be either ${\mathfrak{U}}^{t}_{\zeta}$, ${\mathfrak{U}}^{t}_{\zeta}U^{+}_{\zeta}$, ${\mathfrak{U}}^{b}_{\zeta}$, ${\mathfrak{U}}_{\zeta}$ or $U^{\mathrm{hb}}_{\zeta}$. Then there are algebra homomorphisms $Z_{\mathrm{HC}}\hookrightarrow{\mathfrak{U}}_{\zeta}\rightarrow U$. By the Harish-Chandra isomorphism, we have an algebra homomorphism
${\mathbb{C}}[T/(W,\bullet)]\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda}_{S}).$
By the $S$-linearity, it extends to a homomorphism
@@
$U\text{-}\mathrm{Mod}^{\Lambda}_{S}=\bigoplus_{\omega\in\Xi}U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S},$
such that the Verma module $M^{U}(\lambda)_{S}$ is contained in $U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}$ if and only if $\omega=W_{\mathrm{ex}}\bullet_{l}\lambda$. We have algebra homomorphisms
${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}/W_{\omega}]_{\hat{0}}\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}),\quad\forall\omega\in\Xi.$
The block decomposition induces a decomposition for the subcategory ${\mathcal{O}}^{U}_{S}$, and extends to a decomposition
${\mathcal{O}}^{U}_{R}=\bigoplus_{\omega\in\Xi}{\mathcal{O}}^{U,\omega}_{R},$
for any commutative Noetherian $S$-algebra $R$. Note that when $U={\mathfrak{U}}^{t}_{\zeta}$ or ${\mathfrak{U}}^{b}_{\zeta}$, the category ${\mathcal{O}}$ for $U$ is exactly $U\text{-}\mathrm{mod}^{\Lambda}_{R}$.
@@
Since $u_{\zeta}$, $u_{\zeta}U^{+}_{\zeta}$ are quotients of ${\mathfrak{U}}_{\zeta}^{t}$, ${\mathfrak{U}}^{t}_{\zeta}U^{+}_{\zeta}$, we have block decompositions

### v4

@@ 2.1.1. Affine Weyl group
Let $W_{\mathrm{af}}=W\ltimes{\mathrm{Q}}$ be the affine Weyl group. Let ${\mathrm{I}}_{\mathrm{af}}$ be the subset of simple reflections in $W_{\mathrm{af}}$. Let $W_{\mathrm{ex}}=W\ltimes\Lambda$ be the extended affine Weyl group. We have $W_{\mathrm{ex}}\simeq W_{\mathrm{af}}\rtimes\Lambda/{\mathrm{Q}}$, where $\Lambda/{\mathrm{Q}}$ acts by automorphisms of the affine root system. Denote by $t(\lambda)$ the translation by $\lambda\in\Lambda$ in $W_{\mathrm{ex}}$. For any $n\in{\mathbb{Z}}$, there is an $n$-dilated $(-\rho)$-shifted action of $W_{\mathrm{ex}}$ on $\Lambda$ given by
@@
$\Xi_{\mathrm{sc}}:=\{\lambda\in\Lambda\ |\ 0\leq\langle\lambda+\rho,\check{\alpha}\rangle\leq l,\ \forall\alpha\in\Phi^{+}\}\leftrightarrow\Lambda/(W_{\mathrm{af}},\bullet_{l}).$
@@ 2.2. Quantum groups
We set ${\mathfrak{U}}^{t}_{\zeta}$ be the algebra with triangular decomposition ${\mathfrak{U}}^{t}_{\zeta}=u_{\zeta}^{-}\otimes{\mathfrak{U}}_{\zeta}^{0}\otimes u_{\zeta}^{+}$. For $A=A_{q}$ or $A_{\zeta}$ above, we abbreviate the subalgebras $A^{\leq}:=A^{-}A^{0}$ and $A^{\geq}:=A^{+}A^{0}$.
@@ 2.3. Module categories
The $\Lambda$-action on $\mathscr{U}_{q}^{0}$ preserves the integral forms ${\mathfrak{U}}^{0}_{q}$ and $U^{0}_{q}$, and specializes to an action on ${\mathfrak{U}}^{0}_{\zeta}$ and $U^{0}_{\zeta}$. Let $U^{0}$ be a Noetherian sub-quotient algebra of ${\mathfrak{U}}_{\zeta}^{0}$ or $U_{\zeta}^{0}$ such that the $\Lambda$-action induces an action on $U^{0}$. Let $U=\bigoplus_{\lambda\in{\mathrm{Q}}}U_{\lambda}$ be a Noetherian $\mathrm{Q}$}-graded $\mathbb{C}$}-algebra with triangular decomposition $U=U^{-}\otimes U^{0}\otimes U^{+}$, such that
@@
and satisfies further conditions as in [35, §2.3]. We abbreviate $U^{\leq}=U^{-}U^{0}$ and $U^{\geq}=U^{+}U^{0}$. A deformation ring $R$ for $U$ is a commutative Noetherian $U^{0}$-algebra. Let $\iota:U^{0}\rightarrow R$ be the structure map. If $U^{0}$ is an augmented algebra, then we will view $\mathbb{C}$} as a deformation ring via the counit map $U^{0}\rightarrow{\mathbb{C}}$.
@@
Let $U\text{-}\mathrm{mod}_{R}^{\Lambda}$ be the full subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ consisting of finitely generated $U\otimes R$-modules whose weight spaces are finitely generated $R$-modules. Define the category $\mathcal{O}$} for $U$ to be the full subcategory ${\mathcal{O}}^{U}_{R}$ of $U\text{-}\mathrm{mod}_{R}^{\Lambda}$ of modules that are locally unipotent for the action of $U^{+}$. It is an abelian subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$. If $U^{+}$ is finite dimensional, then we have ${\mathcal{O}}^{U}_{R}=U\text{-}\mathrm{mod}^{\Lambda}_{R}$.
@@
where $R_{\lambda}$ is a $U^{\geq}$-module via $U^{\geq}\twoheadrightarrow U^{0}\xrightarrow{\iota\circ\tau_{\lambda}}R_{\lambda}$. We have an isomorphism $M^{U}(\lambda)_{R}=U^{-}\otimes R_{\lambda}$ as $U^{\leq}$-modules. A Verma flag of an object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ is a filtration of finite length with composition factors by Verma modules. The composition factors of such an object are called Verma factors, which (including multiplicities) are independent on the choice of Verma flags.
@@
The following is standard, see e.g. [23, Prop 3.1].
Lemma 2.1.
We have
(2.3) $\mathrm{Hom}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda\leq\mu,$
(2.4) $\mathrm{Ext}^{1}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda<\mu.$
@@ 2.3.1. Truncated categories
For any subset $\Omega\subset\Lambda$, there is a truncated subcategory (by $\Omega$)
$U\text{-}\mathrm{Mod}_{R}^{\Omega}$
consisting of the module $M$ in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ such that $M_{\mu}=0$ unless $\mu\in\Omega$. It is a Serre subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$. The truncation functor
$\tau^{\Omega}:\ U\text{-}\mathrm{Mod}_{R}^{\Lambda}\rightarrow U\text{-}\mathrm{Mod}_{R}^{\Omega},\quad M\mapsto M\big/U.\big(\bigoplus_{\mu\notin\Omega}M_{\mu}\big)$
by taking the maximal quotient in $U\text{-}\mathrm{Mod}_{R}^{\Omega}$ is left adjoint to the natural inclusion functor $U\text{-}\mathrm{Mod}_{R}^{\Omega}\hookrightarrow U\text{-}\mathrm{Mod}_{R}^{\Lambda}$. We abbreviate $U\text{-}\mathrm{mod}_{R}^{\Omega}:=U\text{-}\mathrm{mod}_{R}^{\Lambda}\cap U\text{-}\mathrm{Mod}_{R}^{\Omega}$.
Let now $\Omega\subset\Lambda$ be a bounded above poset ideal.
Then any finitely generated $U\otimes R$-module in $U\text{-}\mathrm{Mod}_{R}^{\Omega}$ is contained in ${\mathcal{O}}^{U}_{R}$. We set
${\mathcal{O}}^{U,\in\Omega}_{R}:={\mathcal{O}}^{U}_{R}\cap U\text{-}\mathrm{Mod}_{R}^{\Omega}.$
The category ${\mathcal{O}}^{U}_{R}$ is the union of its full subcategories ${\mathcal{O}}^{U,\in\Omega}_{R}$, running over all bounded above poset ideals $\Omega\subset\Lambda$. The category ${\mathcal{O}}^{U,\in\Omega}_{R}$ always admits enough projective objects, in contrast to ${\mathcal{O}}^{U}_{R}$. Moreover, there is a generating family of projective objects in ${\mathcal{O}}^{U,\in\Omega}_{R}$ admitting Verma flags, with factors of the form $M^{U}(\lambda)_{R}$ with $\lambda\in\Omega$, see e.g. [20, Lem 2.3]. If $R$ is local, then any projective module in ${\mathcal{O}}^{U,\in\Omega}_{R}$ admits Verma flags, see e.g. [20, Lem 2.5]. If $R$ is a complete local domain with residue field $\mathbb{F}$}, we write $Q^{U}(\lambda)_{R}^{\Omega}$ as the projective cover of $L^{U}(\lambda)_{\mathbb{F}}$ in ${\mathcal{O}}^{U,\in\Omega}_{R}$, for any $\lambda\in\Omega$.
Let $R^{\prime}$ be a commutative Noetherian $R$-algebra. We have a natural functor
$\otimes_{R}R^{\prime}:\ U\text{-}\mathrm{Mod}^{\Lambda}_{R}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda}_{R^{\prime}}.$
By [20, Prop 2.4], it yields an equivalence
(2.5) $\mathrm{Proj}({\mathcal{O}}^{U,\in\Omega}_{R})\otimes_{R}R^{\prime}{\ \xrightarrow{\sim}\ }\mathrm{Proj}({\mathcal{O}}^{U,\in\Omega}_{R^{\prime}}),$
for any bounded above poset ideal $\Omega$ in $\Lambda$.
@@
Let $\Omega$ be a poset ideal in $\Lambda$. For any object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ admitting a Verma flag, its truncation in $U\text{-}\mathrm{Mod}_{R}^{\Omega}$ is the quotient by the submodule composed by its Verma factors in $\{M^{U}(\lambda)_{R}\}_{\lambda\notin\Omega}$.
Proof.
Let $P$ be an object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ admitting a Verma flag. By (2.4), there is a short exact sequence $0\rightarrow P_{1}\rightarrow P\rightarrow P_{2}\rightarrow 0$, where $P_{1}$ is the submodule composed by the Verma factors in $\{M^{U}(\lambda)\}_{\lambda\notin\Omega}$ and $P_{2}$ is the quotient module composed by the Verma factors in $\{M^{U}(\lambda)\}_{\lambda\in\Omega}$. It is clear that $P_{2}\in U\text{-}\mathrm{Mod}_{R}^{\Omega}$. Note that we have
$\tau^{\Omega}M^{U}(\lambda)_{R}=\begin{cases}M^{U}(\lambda)_{R},&\lambda\in\Omega\\ 0,&\lambda\notin\Omega\end{cases}.$
Using the fact that $\tau^{\Omega}$ is right exact, we deduce that $\tau^{\Omega}(P_{1})=0$ and that $\tau^{\Omega}(P)=\tau^{\Omega}(P_{2})=P_{2}$. ∎
For $\Omega=\{\lambda\in\Lambda|\lambda\leq\nu\}$, we abbreviate $\Omega$” and $\in\Omega$” by $\leq\nu$” in all the superscripts above. We will drop the superscripts $U$ from the notations above when $U=U^{\mathrm{hb}}_{\zeta}$. For example, if $\Omega=\{\lambda\in\Lambda|\lambda\leq\nu\}$ and $U=U^{\mathrm{hb}}_{\zeta}$, we abbreviate ${\mathcal{O}}^{U,\in\Omega}_{R}$, $Q^{U}(\lambda)_{R}^{\Omega}$ by ${\mathcal{O}}^{\leq\nu}_{R}$, $Q(\lambda)_{R}^{\leq\nu}$.
@@ 2.3.2. Projectives and simples in ${\mathcal{O}}_{R}$
We set
@@
Then $S$ is naturally an algebra of ${\mathfrak{U}}^{0}_{\zeta}={\mathbb{C}}\langle K_{\lambda}\rangle_{\lambda\in\Lambda}={\mathbb{C}}[T]$. We thus view $S$ as a deformation ring for both ${\mathfrak{U}}^{b}_{\zeta}$ and $U^{\mathrm{hb}}_{\zeta}$.
Lemma 2.3 ([35, Lem 3.1]).
@@
    Let $\Omega$ be a poset ideal in $\Lambda$, and let $\lambda\in\Omega$. The projective cover for $L(\lambda)_{\mathbb{C}}$ in ${\mathcal{O}}^{\in\Omega}_{S}$ (resp. ${\mathcal{O}}^{\in\Omega}_{{\mathbb{C}}}$) is $Q(\lambda)_{S}^{\Omega}=\tau^{\Omega}Q(\lambda)_{S}$ (resp. $Q(\lambda)_{\mathbb{C}}^{\Omega}=\tau^{\Omega}Q(\lambda)_{\mathbb{C}}$).
@@ 2.3.3. Block decomposition
Till the end of this section, we let $U$ be either ${\mathfrak{U}}^{t}_{\zeta}$, ${\mathfrak{U}}^{t}_{\zeta}U^{+}_{\zeta}$, ${\mathfrak{U}}^{b}_{\zeta}$, ${\mathfrak{U}}_{\zeta}$ or $U^{\mathrm{hb}}_{\zeta}$. Then there are algebra homomorphisms $Z_{\mathrm{HC}}\hookrightarrow{\mathfrak{U}}_{\zeta}\rightarrow U$. Let $R$ be a commutative Noetherian $S$-algebra. We have algebra homomorphisms
$S\otimes Z_{\mathrm{HC}}\rightarrow R\otimes Z_{\mathrm{HC}}\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda}_{R}).$
By the Harish-Chandra isomorphism, it factors through a homomorphism
@@
$U\text{-}\mathrm{Mod}^{\Lambda}_{R}=\bigoplus_{\omega\in\Xi}U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{R},$
such that the Verma module $M^{U}(\lambda)_{R}$ is contained in $U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{R}$ if and only if $\omega=W_{\mathrm{ex}}\bullet_{l}\lambda$. We have algebra homomorphisms
(2.8) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}/W_{\omega}]_{\hat{0}}\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}),\quad\forall\omega\in\Xi.$
The block decomposition induces a decomposition
${\mathcal{O}}^{U}_{R}=\bigoplus_{\omega\in\Xi}{\mathcal{O}}^{U,\omega}_{R}.$
Since $u_{\zeta}$, $u_{\zeta}U^{+}_{\zeta}$ are quotients of ${\mathfrak{U}}_{\zeta}^{t}$, ${\mathfrak{U}}^{t}_{\zeta}U^{+}_{\zeta}$, we obtain block decompositions
@@
For a subset $\Omega\subset\Lambda$, we abbreviate $U\text{-}\mathrm{Mod}^{\Omega,\omega}_{S}:=U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}\cap U\text{-}\mathrm{Mod}^{\Omega}_{S}$ and ${\mathcal{O}}^{U,\omega,\in\Omega}_{R}:={\mathcal{O}}^{U,\omega}_{R}\cap U\text{-}\mathrm{Mod}^{\Omega}_{S}$. Again, for $\Omega=\{\lambda\in\Lambda|\lambda\leq\nu\}$, we abbreviate $\Omega$” and $\in\Omega$” by $\leq\nu$” in the superscripts.
We will also drop the superscripts $U$ from the notations above when $U=U^{\mathrm{hb}}_{\zeta}$.